OpenStudy (anonymous):

If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? @Mathematics

OpenStudy (anonymous):

what is the formula for volume of sphere??????

OpenStudy (anonymous):

see how u will solve it on ur own

OpenStudy (anonymous):

Volume = 4/3 * pi * r^3 Surface Area = 4 * pi * r^2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

we have to calculate the radius of the sphere because that is only changing and other terms are constant......right

OpenStudy (anonymous):

yes.. i am stuck since 2-3 days

OpenStudy (anonymous):

so vikrant hvn any idea how to calculate the radius....?

OpenStudy (anonymous):

nup :(

OpenStudy (anonymous):

give a try dont give up

OpenStudy (anonymous):

ok will give u hint let us suppose b4 volume was 100 now the volume is 172.8 so u can solve for radius

OpenStudy (anonymous):

so for solving for r u will get 2.87 and 3.45 then u can easily calculate the value for surface area increase

OpenStudy (anonymous):

in this case answer is approximately 44.5 %

OpenStudy (anonymous):

I am with sheg!

OpenStudy (anonymous):

calculate and see

OpenStudy (anonymous):

U said 172.8 = 4/3 pi r^3 CUBERT(129.6/pi) = r How can i calculate a cube root on a paper ????????